# A Security Risk Assessment Method for Distributed Ledger Technology (DLT) based Applications: Three Industry Case Studies

### v1

@@ III-C Research Method
Double Spending Attack
When the same single digital tokens can be spent on more than one because the attacker manages to spend both of them simultaneously. Double spending is one of the major security issues in most blockchain systems, but it is difficult to launch unless an adversary has massive computing power successfully.
Network
Network Layer
Finney attack; race attacks; 51 % attacks
[35]
Eclipse Attack
a set of malicious, colluding overlay nodes arranges for a correct node to peer only with coalition members. If successful, the attacker can mediate most or all communication to and from the victim. By supplying biased neighbor information during normal overlay maintenance, a modest number of malicious nodes can eclipse a large number of correct victim nodes.
Network
Network Layer
Consensus Majority Attack, Engineering block races, Splitting mining power (51% attack), selfish mining, 0-confirmation double spend, N-confirmation double-spend
Sybil attack
[36]
Long Range Attack
History Attack
Proof of Stake vulnerability, due to weak subjectivity and costless simulation, an attacker attempts to rewrite the history of a ledger by forking the ledger
Transaction information
Consensus Layer
[37]
@@ IV Systematic literature review
[42]
2002
Literature study
Information security
Qualitative
NIST
[43]
2002
Literature study
Case study
Qualitative
[44]
2009
Literature study
Qualitative
ISO 27001
[45]
2014
Literature study
Experiment
Cloud computing
Qualitative
ISO 27005
[46]
2014
Literature study, Benchmarking
Benchmarking
Smart grid
Qualitative
CORAS, ISO 27001
[47]
2015
Literature study, Benchmarking
Benchmarking
Medical devices
Qualitative
ISO 27001
[48]
2016
Literature study, Benchmarking
Benchmarking
Information system security
Qualitative
ISO 27002
[49]
2016
Literature study, Benchmarking
Benchmarking
SCADA systems
Qualitative
ISO 31000, NIST SP 800-30
[50]
2018
Document Analysis, Literature study
Benchmarking
Blockchain
Qualitative
NIST SP-800-30
[51]
2018
Design science, Literature study
Case study
DLT applications
Qualitative
ISO 27005
[52]
2018
Literature study
Information
Qualitative
ISO 27005, NIST SP-800-30
[53]
2018
Literature study
IoT
Quantitative
ISO 31000, NIST SP 800-30
[54]
2019
Design science, Literature study
Cloud
Qualitative
ISO 31000, NIST SP 800-30
[55]
2021
Document Analysis, Literature study
Blockchain
Quantitative
ISO 27005, ISO 31000
[56]
2021
Document Analysis, Literature study
Automotive Cybersecurity
Quantitative
ISO 26262
[57]
2021
Fuzzy set theory, Literature study
Experiment
Control System
Quantitative
ISO 26262
[58]
2021
Literature study, Benchmarking
Benchmarking
Quantitative
[59]
2021
Literature study, Benchmarking
Benchmarking
Funding
Quantitative
[60]
2022
Literature study, Benchmarking, Case study
Case study
Information security
Qualitative
ISO/IEC 27005, CORAS
[61]
2022
Document Analysis, Literature study , Benchmarking
Benchmarking
Cloud
Qualitative
NIST SP 800:30, ISO 27001, SO 30001
[62]
2022
Literature study, Benchmarking, Case study
Case study
Quantitative
SO 27001, NIST SP 800:30
[63]
2022
Literature study
Distributed Ledger
Qualitative
ISO 27005, ISO 27001
[64]
2022
Literature study, Benchmarking
Benchmarking
Blockchain
Qualitative
[65]
2022
Document Analysis, Literature study
Benchmarking
SCADA system
Quantitative
[66]
2022
Document Analysis, Literature study , Benchmarking
Benchmarking
IT infrastructure
Qualitative
@@ VI-A SecureSECO
In this section, we outline all the steps of SRAMDA that were conducted for the SearchSECO project.Figure IV shows the summary of employing these steps.
@@ VI-A1 Gather project specifications
SecureSECO aims to secure and increase trust in the software ecosystem through DLT and empirical software engineering research. SearchSECO aim to create a secure software ecosystem. This project uses DLT to store valuable trust facts about software and users. This project relies on trust, which is hard to measure, but it is possible to protect and use smart contracts and consensus mechanisms. The project leaders determined the scope of risk assessment during the interview as follows: To provide valuable insights and awareness of cybersecurity risks. It can provide the assessment results for further designing and developing the project.
TABLE IV: A step-by-step application of SRAMDA to SearchSECO.

### v2

@@ III-C Research Method
Double Spending Attack - When the same single digital tokens can be spent on more than one because the attacker manages to spend both of them simultaneously. Double spending is one of the major security issues in most blockchain systems, but it is difficult to launch unless an adversary has massive computing power successfully. Network Network Layer - Finney attack; race attacks; 51 % attacks [35]
Eclipse Attack - a set of malicious, colluding overlay nodes arranges for a correct node to peer only with coalition members. If successful, the attacker can mediate most or all communication to and from the victim. By supplying biased neighbor information during normal overlay maintenance, a modest number of malicious nodes can eclipse a large number of correct victim nodes. Network Network Layer Consensus Majority Attack, Engineering block races, Splitting mining power (51% attack), selfish mining, 0-confirmation double spend, N-confirmation double-spend Sybil attack [36]
Long Range Attack History Attack Proof of Stake vulnerability, due to weak subjectivity and costless simulation, an attacker attempts to rewrite the history of a ledger by forking the ledger Transaction information Consensus Layer - - [37]
@@ IV Systematic literature review
[42] 2002 Literature study - Information security Qualitative NIST
[43] 2002 Literature study Case study - Qualitative -
[44] 2009 Literature study - - Qualitative ISO 27001
[45] 2014 Literature study Experiment Cloud computing Qualitative ISO 27005
[46] 2014 Literature study, Benchmarking Benchmarking Smart grid Qualitative CORAS, ISO 27001
[47] 2015 Literature study, Benchmarking Benchmarking Medical devices Qualitative ISO 27001
[48] 2016 Literature study, Benchmarking Benchmarking Information system security Qualitative ISO 27002
[49] 2016 Literature study, Benchmarking Benchmarking SCADA systems Qualitative ISO 31000, NIST SP 800-30
[50] 2018 Document Analysis, Literature study Benchmarking Blockchain Qualitative NIST SP-800-30
[51] 2018 Design science, Literature study Case study DLT applications Qualitative ISO 27005
[52] 2018 Literature study - Information Qualitative ISO 27005, NIST SP-800-30
[53] 2018 Literature study - IoT Quantitative ISO 31000, NIST SP 800-30
[54] 2019 Design science, Literature study - Cloud Qualitative ISO 31000, NIST SP 800-30
[55] 2021 Document Analysis, Literature study - Blockchain Quantitative ISO 27005, ISO 31000
[56] 2021 Document Analysis, Literature study - Automotive Cybersecurity Quantitative ISO 26262
[57] 2021 Fuzzy set theory, Literature study Experiment Control System Quantitative ISO 26262
[58] 2021 Literature study, Benchmarking Benchmarking - Quantitative -
[59] 2021 Literature study, Benchmarking Benchmarking Funding Quantitative -
[60] 2022 Literature study, Benchmarking, Case study Case study Information security Qualitative ISO/IEC 27005, CORAS
[61] 2022 Document Analysis, Literature study , Benchmarking Benchmarking Cloud Qualitative NIST SP 800:30, ISO 27001, SO 30001
[62] 2022 Literature study, Benchmarking, Case study Case study - Quantitative SO 27001, NIST SP 800:30
[63] 2022 Literature study - Distributed Ledger Qualitative ISO 27005, ISO 27001
[64] 2022 Literature study, Benchmarking Benchmarking Blockchain Qualitative -
[65] 2022 Document Analysis, Literature study Benchmarking SCADA system Quantitative -
[66] 2022 Document Analysis, Literature study , Benchmarking Benchmarking IT infrastructure Qualitative -
@@ VI-A SecureSECO
In this section, we outline all the steps of SRAMDA that were conducted for the SecureSECO project.Figure IV shows the summary of employing these steps.
@@ VI-A1 Gather project specifications
SecureSECO aims to secure and increase trust in the software ecosystem through DLT and empirical software engineering research. SecureSECO aim to create a secure software ecosystem. This project uses DLT to store valuable trust facts about software and users. This project relies on trust, which is hard to measure, but it is possible to protect and use smart contracts and consensus mechanisms. The project leaders determined the scope of risk assessment during the interview as follows: To provide valuable insights and awareness of cybersecurity risks. It can provide the assessment results for further designing and developing the project.
TABLE IV: A step-by-step application of SRAMDA to SecureSECO.